If the period of an oscillation is 0.0002 ms

what is the period in seconds?
Answer #1

period of an oscillation T = 0.0002 ms

1 ms = 10-3 s

T = 0.0002 * 10-3

= 0.0000002 s

period in seconds = 2 * 10-7 s